The maximum torque experienced by a coil in a 8.00 × 10-3 T magnetic field is 0.70 N.m. The coil is circular and consists of three turns. The current in the coil is 9.10 A. What is the length of the wire from which the coil is made?
Answer #1

maximum torque n the coil= T= N I A B

A= area of coil,

0.70 = 3 (9.1)(A)(8x10^-3)

A=3.205 m^2

pi r^2 = 2.205

r= 1.01 m

length of wire= number of turns * circumfrence of circle

L= 3 (2 pi r)

L= 3 (2* pi * 1.01)

L=19.03 m